OpenStudy (anonymous):

Solve for x. 5+2x-10+12/3=25 @One098

OpenStudy (one098):

\[5+2x-10+\frac{ 12 }{ 3 }=25\]First off, with the fraction, divide 12 by 3. \[5+2x-10+4=25\]Your result was 4. Now subtract 4 from both sides. \[5+2x-10+4-4=25-4\]The two 4's cancel out, then simply subtract the right side. \[5+2x-10=21\]Now add 10 to both sides. \[5+2x-10+10=21+10\]The two 10's cancel out and just add the right side. \[5+2x=31\]Now subtract 5 from both sides. \[5-5+2x=31-5\]The two 5's cancel out and subtract the left side. \[2x=26\]Now divide 2 by both sides. \[\frac{ 2x }{ 2 }=\frac{ 26 }{ 2 }\]The two 2's cancel out and divide the right side. \[x=13\]To check to make sure this is your answer, plug it in. \[5+2\times13-10+\frac{ 12 }{ 3 }=25\]\[5+26-10+4=25\]\[31-10+4=25\]\[21+4=25\]\[25=25\] So x=13.

OpenStudy (anonymous):

ohh thank you so much that helped alot